Researchers at a pharmaceutical company have found that the effective time duration of a safe dosage of a pain relief drug is normally distributed with mean 2 hours and standard deviation 0.3 hour. For a patient selected at random:
a) What is the probability that the drug will be effective for 2 hours or less?
b) What is the probability that the drug will be effective for 1 hour or less?
c) What is the probability that the drug will be effective for 3 hours or more?

Answer by stanbon(48568)

Researchers at a pharmaceutical company have found that the effective time duration of a safe dosage of a pain relief drug is normally distributed with mean 2 hours and standard deviation 0.3 hour. For a patient selected at random:
a) What is the probability that the drug will be effective for 2 hours or less?
That is the left half of the area under the normal curve; P(x<=2 hrx)= 1/2
---------------------------
b) What is the probability that the drug will be effective for 1 hour or less?
t(1) = (1-2)/0.3 = -1/0.3 = -10/3
P(-10< t < -10/3 for df= large) = 0.000602
------------------------------
c) What is the probability that the drug will be effective for 3 hours or more?
t(3) = (3-2)/0.3 = 1/0.3 = 10/3
P(10/3 < t < 10 for df= large) = 0.000602
==============
Cheers,
Stan H.